Jose is 12 years older than Maria. The sum of their ages is 40. How old is Maria? How old is Jose?
Answer by jndarrell(58)

Maria = x
Jose = x+12
Sum (addition) total of two is 40
x+x+12=40
Simplify
2x+12=40
Subtract 12 from both sides
2x=28
Divide both sides to get x alone
x=14
Plug answer into original people
Maria = 14yrs
Jose = 26yrs